OpenStudy (anonymous):

Which statement is true about the product: √2( ³√2+ √18)? A. it is rational and equal to 12 B.it is rational and equal to 18 C. It is irrational and equal to 6 + 2 √18 D. It is irrational and equal to 3 + √18

OpenStudy (narissa):

well, if the question is 2(3√2+18) then 6√2+36 is the correct answer

OpenStudy (anonymous):

Oh okay, Um thank you but maybe I wrote that down wrong because these are the only answer choices they gave me.

OpenStudy (narissa):

yeah the answer choices dont really make sence but this is how you solve it a(b+c)=ab+ac √2(3 √2+ √18)=(2x3√2) + (2x18)= 6√2+36

OpenStudy (anonymous):

Thank you again for helping me out anyways!! I guess I can always just email my teacher and ask her how to solve it!!

OpenStudy (narissa):

oh okay r u in k12

OpenStudy (anonymous):

No, I'm not

OpenStudy (narissa):

oh.. i can tag someone to help u if u like

OpenStudy (narissa):

@ganeshie8

OpenStudy (anonymous):

is it the square root of 2 times the cube root of 2 plus the square root of 18?

OpenStudy (anonymous):

Im having a hard time making it out

OpenStudy (anonymous):

square root of 2( ³ square root of 2 + square root of 18 : This is the best way I can explain it

OpenStudy (anonymous):

is that 3 times the square root of 2?

OpenStudy (anonymous):

and not the cube root of two

OpenStudy (anonymous):

that 3 looks like an index its so small

OpenStudy (anonymous):

Oh, I'm so sorry!! I guess so its kind of hard to understand its off of Florida virtual school they don't give very much of an explanation! But it's the cubed root of two squared + square root of 18 unless I'm misinterpreting it and it is 3 times the square root of 2

OpenStudy (anonymous):

can you take a screenshot of the original expression

OpenStudy (anonymous):

No I can't I tried but it won't let me I'm sorry for the confusion I will just have to ask my teacher for help but thank you for trying!!

OpenStudy (anonymous):

@Brian42